Michael is a bicycle commuter. he has observed that the reduces the time for his 9 km commute by 15 min when he increases his average speed by 3 km/h. what is Michael's faster speed?


1
Expert's answer
2021-09-12T23:52:28-0400

Let "x=" Michael's slower speed. Then his faster speed will be "x+3."

When Michael travels with slower speed his time is "t_1=\dfrac{9}{x}."

When Michael travels with faster speed his time is "t_2=\dfrac{9}{x+3}."

Michael reduces the time for his 9 km commute by 15 min


"t_1-t_2=\dfrac{1}{4}"

"\dfrac{9}{x}-\dfrac{9}{x+3}=\dfrac{1}{4}"

"36(x+3)-36x=x(x+3)"

"x^2+3x-108=0, x>0"

"D=(3)^2-4(1)(108)=441"

"x=\dfrac{-3\pm\sqrt{441}}{2(1)}=\dfrac{-3\pm21}{2}"

Since "x>0," we take


"x=\dfrac{-3+21}{2}"

"x=9"

Therefore Michael's faster speed is "12" km/h.